# On the torsion growth in quadratic number fields for elliptic curves defined over the rationals

### v1

@@ 4. The prime $\ell=3$
$\begin{cases}x=\displaystyle\frac{x^{\prime}}{4}\\ \\ y=\displaystyle\frac{y^{\prime}}{8}-\frac{1}{2}(a_{1}\frac{x}{4}+a_{3})\end{cases}$
@@
and we can use the fact([2, Corollary 4]) that if $n>1$ is an odd integer, then
$E(K)[n]=E(\mathbb{Q})[n]\times E_{d}(\mathbb{Q})[n].$
@@
$x^{\prime}=dx,\quad y^{\prime}=d\sqrt{d}y$
@@
$E_{4}:y^{2}+\overline{a}_{1}xy+\overline{a}_{3}y=x^{3}+\overline{a}_{2}x^{2}$
@@
$\left\{\begin{array}[]{lcl}s&=&\displaystyle\frac{8d^{2}b_{4}+2db_{2}x_{1}+3x_{1}^{2}}{2y_{1}}\\ \overline{a}_{1}&=&2s\\ \overline{a}_{2}&=&-s^{2}+3x_{1}+db_{2}\\ \overline{a}_{3}&=&2y_{1}\\ \Delta_{4}&=&2^{12}d^{6}\Delta_{1}\end{array}\right.$
We have that $\overline{a}_{2}\not=0$ because $P$ does not have order $3$ (see [7, V.5]). So we can make a change of variables to get the Tate normal form
@@
Using [4, Example 4.6], there exists $t\in\mathbb{Q}$ such that
$\left\{\begin{array}[]{lcl}b&=&\displaystyle(t-1)t^{2}(t^{2}-t+1)\\ c&=&(t-1)t^{2}\\ \Delta_{b,c}&=&(t-1)^{9}t^{9}(t^{2}-t+1)^{3}(t^{3}-6t^{2}+3t+1)\end{array}\right.$
@@
Because of the Nagell-Lutz theorem (as explained in the subsection for the cases $N=8,16$), $y_{1}^{2}|d^{6}2^{12}\Delta_{1}$, which implies that $v_{3}(y_{1})\in\{0,1,2,3\}$ because $d$ is squarefree. For each of these values, the last equation gives us a contradiction.
@@
Therefore, $v_{3}(y_{1})\in\{0,1\}$. On the other hand, we have the relations $-b\overline{a}_{3}^{2}=\overline{a}_{2}^{3}$ and $c=1-(\overline{a}_{1}\overline{a}_{2})/\overline{a}_{3}$, with
$\begin{array}[]{lcl}b&=&\displaystyle\frac{(p-q)p^{2}(p^{2}-pq+q^{2})}{q^{5}}\\ c&=&\displaystyle\frac{(p-q)p^{2}}{q^{3}}.\end{array}$
@@
First, we show that $v_{3}(p-q)=0$. Let us assume $v_{3}(p-q)>0$, we have that
$v_{3}(p^{2}-pq+q^{2})=v_{3}((p-q)^{2}+pq)=0$
and as $0<3v_{3}(p-q)=v_{3}(p^{3}-q^{3}-3pq(p-q))$, $v_{3}(p^{3}-q^{3})>0$.
Therefore
$v_{3}(p^{3}+q^{3}-6p^{2}q+3pq^{2})=v_{3}(p^{3}-q^{3}+2q^{3}-6p^{2}q+3pq^{2})=0.$
So, Equation (21) becomes
$6v_{3}(d)=4v_{3}(y_{1})+5v_{3}(p-q).$
We know that $v_{3}(d)\in\{0,1\}$ (because $d$ is squarefree) and in both cases we get a contradiction. So we have that $v_{3}(p-q)=0$. Now, $v_{3}(p-q)=0$ implies that $v_{3}(p+q)>0$, because the only possibilities are $p\equiv 1\mod 3$ and $q\equiv-1\mod 3$ or vice versa. It implies that
$v_{3}(p^{3}+q^{3}-6p^{2}q+3pq^{2})>0.$
Moreover, $p=\pm 1+3k$ and $q=\mp 1+3r$. Making the substitution in the equation above we get
$27k^{3}-162k^{2}r+81kr^{2}+27r^{3}\pm 81k^{2}\mp 162kr+54k-27r\pm 9,$
which implies that $v_{3}(p^{3}+q^{3}-6p^{2}q+3pq^{2})=2$. Similarly, making the substitution in $p^{2}-pq+q^{2}$, we get
$39k^{2}-9kr+9r^{2}\pm 9k\mp 9r+$ 3 .
So $v_{3}(p^{2}-pq+q^{2})=1$, and Equation (21) becomes
@@
If $E(\mathbb{Q})[3]=\mathcal{C}_{3}$, the hypothesis implies that $E_{\mathrm{tors}}(K)=E_{\mathrm{tors}}(\mathbb{Q})\times\mathcal{C}_{3}$, because of Theorem 2 of [2]. Therefore we can assume that $E(\mathbb{Q})[3]$ is trivial.
@@
Now we go by contradiction. Using Theorem $2$ of [2], we see that the possible cases that can arise are:
@@
  • $E_{\mathrm{tors}}(\mathbb{Q})=\mathcal{C}_{4}$ and $E_{\mathrm{tors}}(K)=\mathcal{C}_{2}\times\mathcal{C}_{12}$, or else
@@
Case I: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$, $\mathcal{C}_{4}\leq E_{\mathrm{tors}}(K)$. As in subsection 3 we can do a change of variables and we get a Weierstrass equation:
@@
Now Lemma 1 tell us that $K=\mathbb{Q}(\sqrt{A\pm 2s})$ with $s\in\mathbb{Q}$ such that $s^{2}=B$, so $s\in\mathbb{Z}$. Let us call $v_{3}$ the $3$-adic valuation. As $3|d$, $A\pm 2s$ must be a multiple of $3$. Then, from the equation relating the discriminants, we obtain
@@
Case II: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$, $\mathcal{C}_{2}\times\mathcal{C}_{2}\leq E_{\mathrm{tors}}(K)$. We can compute, as in Section 2, the $2$-torsion points with the $2$-division polynomial $\psi_{2}=2y+a_{1}x+a_{3}$; substituting
@@
we obtain the expression (see, again [11, Ch. III])
@@
[…]
@@
Case III: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}\times\mathcal{C}_{2}$, $\mathcal{C}_{4}\times\mathcal{C}_{2}\leq E_{\mathrm{tors}}(K)$. Again, as in Subsection 3, we can do a change of variables and we get a Weierstrass equation:
@@
In this situation, like in Subsection 3, we can assume that there is a point $Q\in E(K)[4]$ such that $2Q=(0,0)$. Applying Lemma 2 as in Subsection 3, we can assume that the polynomial $x^{3}+Ax^{2}+Bx$ factors as $x(x-\alpha)(x-\beta)$, where $\alpha,\beta\in\mathbb{Q}$ satisfy that $-\alpha$ and $-\beta$ are squares in $K^{*}$, but they are not both squares in $\mathbb{Q}^{*}$. Thus either $-\alpha=a^{2}$ and $\-\beta=db^{2}$, or $-\alpha=da^{2}$ and $\-\beta=b^{2}$, or else $-\alpha=da^{2}$ and $\-\beta=db^{2}$ for some rational integers $a$ and $b$. Replacing $B=\alpha\beta$ and $A=-\alpha-\beta$ in the equation for $\Delta$ above, we obtain in all cases that there is an integer $V\in\mathbb{Z}$ such that
@@ 5. The primes $\ell=5$ and $\ell=7$
Throught this section we can already assume that the curve $E$ does not have complex multiplication, as the CM case has already been solved by E. González–Jiménez [1].
@@
Next, we will see the form of $I(\Lambda/\ell)$ according to the type of reduction of $E$.
5.1. Ordinary reduction with nontrivial wild inertia action
The following proposition, which is proven in [10, Section 1.11], exactly determines the image of $I_{\ell}$ under $\overline{\rho}_{E,\ell}$:
Proposition 9 ([10]).
Let $E/\mathbb{Q}$ be an elliptic curve with good reduction of height 1. Then one and only one of the following possibilities holds:
@@ 5.1. Ordinary reduction with nontrivial wild inertia action
    The wild inertia group $I^{\mathrm{wild}}_{\ell}$ does not act trivially on $E[\ell]$. Then the image of $I^{\mathrm{wild}}_{\ell}$ is a cyclic group of order $\ell$, and in a suitable basis, it can be represented as
@@
We will assume in this section we are in the second case, where $I^{\mathrm{wild}}_{\ell}$ does not act trivially. Note that $H_{3}$ does not leave any element other than $(0,0)^{t}$ invariant; in particular, there cannot be any rational point of $\ell$-torsion.
Furthermore, no subgroup of index $2$ in $H_{3}$ leaves any element other than $(0,0)^{t}$ invariant. Indeed, any subgroup of index $2$ must contain the subgroup $H_{2}$. On the other hand, we have
@@
if and only if $by=0$, so the only nontrivial fixed points by the entire subgroup are of the form $(x,0)^{t}$, where $x\neq 0$. But any subgroup of index $2$ must contain at least one matrix of the form
@@
and this matrix does not fix the point $(x,0)^{t}$. This shows that for $\ell=5,7$, it is not possible to add torsion in an extension that ramifies at $\ell$.
We will address now the case where the inertia group $I^{\mathrm{wild}}_{\ell}$ acts trivially.
5.2. Ordinary Reduction with Trivial Wild Inertia Action
@@
$\ell=5$. In this case, [14] shows that $\overline{\rho}_{E,5}(G_{\mathbb{Q}})$ is conjugate in $\mathop{\mathrm{GL}}\nolimits_{2}(\mathbb{F}_{5})$ to a group from a list of 15 possible groups (see [14, Theorem 1.4]). Furthermore, Sutherland studies these groups and determines in each case the index of the largest subgroup that fixes a nonzero vector in $\mathbb{F}_{5}^{2}$; this quantity coincides with the degree of the minimal extension $K/\mathbb{Q}$ such that $E$ has a rational point of $5$-torsion.
In Table 3 (page 64 of [12]), we can find the list of these 15 groups along with their indices. There are only four of them where this degree is $2$, specifically those labeled as 5Cs.1.3, 5Cs.4.1, 5B.1.4, and 5B.4.1. Let us examine each of these cases:
  • 5Cs.1.3. It is a cyclic subgroup of order $4$, generated by the matrix $\begin{pmatrix}3&0\\ 0&4\end{pmatrix}$. Since we are assuming that $E$ has good ordinary reduction at $5$, $\overline{\rho}_{E,\ell}(G_{\mathbb{Q}})$ must contain a subgroup conjugate to $H_{1}$, which is also cyclic of order $4$. However, the subgroup 5Cs.1.3 is not conjugate to $H_{1}$, because it does not fix any nontzero element of $\mathbb{F}_{5}^{2}$. The conclusion is that this image cannot occur if $E$ has good ordinary reduction at $5$ (it also cannot have supersingular reduction, by a similar reasoning).
@@
    The subgroup generated by the matrix $\begin{pmatrix}1&0\\ 0&2\end{pmatrix}$ is conjugate to the subgroup $H_{1}$, and it is the only cyclic subgroup of order $4$ in 5Cs.4.1 that fixes a nonzero element of $\mathbb{F}_{5}^{2}$. Therefore,
@@
    Moreover, $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$ is a subgroup of $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/\mathbb{Q})$ of order $4$. Since $L\neq K$, it must a subgroup of $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/\mathbb{Q})$ different from $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)$, and thus it does cannot fix a nonzero element of $\mathbb{F}_{5}^{2}$.
@@
    Again, there is a subgroup which is conjugated to $H_{1}$, the subgroup
    $\left\{\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
@@
    In this case there are 4 more cyclic subgroups of order $4$ that leave a fixed point; specifically:
    $\left\{\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\begin{pmatrix}1&3\\ 0&4\end{pmatrix},\begin{pmatrix}1&2\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\},\quad\left\{\begin{pmatrix}1&1\\ 0&3\end{pmatrix},\begin{pmatrix}1&4\\ 0&4\end{pmatrix},\begin{pmatrix}1&3\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&3\end{pmatrix}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\},\quad\left\{\begin{pmatrix}1&3\\ 0&3\end{pmatrix},\begin{pmatrix}1&2\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    In fact, all of these groups are conjugated to $H_{1}$:
    $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}^{-1}=\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\quad\begin{pmatrix}1&1\\ 0&2\end{pmatrix}\begin{pmatrix}1&0\\ 0&3\end{pmatrix}\begin{pmatrix}1&1\\ 0&2\end{pmatrix}^{-1}=\begin{pmatrix}1&1\\ 0&3\end{pmatrix}$
    $\begin{pmatrix}1&2\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\begin{pmatrix}1&2\\ 0&1\end{pmatrix}^{-1}=\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\quad\begin{pmatrix}1&3\\ 0&2\end{pmatrix}\begin{pmatrix}1&0\\ 0&3\end{pmatrix}\begin{pmatrix}1&3\\ 0&2\end{pmatrix}^{-1}=\begin{pmatrix}1&3\\ 0&3\end{pmatrix}$
    By choosing a basis, we can assume that $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)=H_{1}$.
@@
    If there exists a point of $5$-torsion over $K$, the only possibility is that the group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$, which is a group of order $20$, is the union of all elements whose upper left entry is $1$. This can be seen by looking at the list of elements of the group 5B.4.1; the only 20 elements that fix the same element of $\mathbb{F}_{5}^{2}$ are these.
@@
    But in that case, $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)=H_{1}\subset\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$, and therefore $K\subset L$. Since we have seen that $K$ and $L$ are linearly disjoint over $\mathbb{Q}$, this is a contradiction.

### v2

@@ 4. The prime $\ell=3$
$\begin{cases}x=\displaystyle\frac{x^{\prime}}{4},\\ \\ y=\displaystyle\frac{y^{\prime}}{8}-\frac{1}{2}(a_{1}\frac{x}{4}+a_{3}),\end{cases}$
@@
and we can use the fact ([3, Corollary 4]) that if $n>1$ is an odd integer, then
$E(K)[n]\cong E(\mathbb{Q})[n]\times E_{d}(\mathbb{Q})[n].$
@@
$x^{\prime}=dx,\quad y^{\prime}=d\sqrt{d}y,$
@@
$E_{4}:y^{2}+\overline{a}_{1}xy+\overline{a}_{3}y=x^{3}+\overline{a}_{2}x^{2},$
@@
$\left\{\begin{array}[]{lcl}s&=&\displaystyle\frac{8d^{2}b_{4}+2db_{2}x_{1}+3x_{1}^{2}}{2y_{1}}\\ \overline{a}_{1}&=&2s,\\ \overline{a}_{2}&=&-s^{2}+3x_{1}+db_{2},\\ \overline{a}_{3}&=&2y_{1},\\ \Delta_{4}&=&2^{12}d^{6}\Delta_{1}.\end{array}\right$.
We have that $\overline{a}_{2}\not=0$ because $P$ does not have order $3$ (see [8, V.5]). So we can make a change of variables to get the Tate normal form
@@
Using [5, Example 4.6], there exists $t\in\mathbb{Q}$ such that
$\left\{\begin{array}[]{lcl}b&=&\displaystyle(t-1)t^{2}(t^{2}-t+1),\\ c&=&(t-1)t^{2},\\ \Delta_{b,c}&=&(t-1)^{9}t^{9}(t^{2}-t+1)^{3}(t^{3}-6t^{2}+3t+1).\end{array}\right$.
@@
Because of the Nagell-Lutz theorem (as explained in the subsection for the cases $N=8,16$), $y_{1}^{2}|d^{6}2^{12}\Delta_{1}$, which implies that $v_{3}(y_{1})\in\{0,1,2,3\}$ because $d$ is squarefree. For each of these values, the last equation gives us a contradiction, as the left side is strictly greater than the right side.
@@
Therefore, $v_{3}(y_{1})\in\{0,1\}$. On the other hand, we have the relations $b\overline{a}_{3}^{2}=\overline{a}_{2}^{3}$ and $c=1-(\overline{a}_{1}\overline{a}_{2})/\overline{a}_{3}$, with
$b=\frac{(p-q)p^{2}(p^{2}-pq+q^{2})}{q^{5}},\qquad c=\frac{(p-q)p^{2}}{q^{3}}.$
@@
We can rewrite (21) in terms of $t=p/q$ and we get
(22) $6v_{3}(d)+v_{3}(t^{2}-t+1)=4v_{3}(y_{1})+5v_{3}(t-1)+v_{3}(t^{3}-6t^{2}+3t+1).$
As $v_{3}(t)=0$, $t\equiv\pm 1\mod 3$. Let us assume first that $t=1+3k$ for some $k\in\mathbb{Q}$ with $v_{3}(k)\geq 0$. Substituting $t$, we get
$t^{2}-t+1$ $=$ $9k^{2}+3k+1,$
$t^{3}-6t^{2}+3t+1$ $=$ $343k^{3}+147k^{2}-42k-17.$
which implies that $v_{3}(t^{2}-t+1)=v_{3}(t^{3}-6t^{2}+3t+1)=0$. Therefore, (22) becomes
$6v_{3}(d)=4v_{3}(y_{1})+5v_{3}(t-1).$
Using that $v_{3}(t-1)>0$ and $v_{3}(d)\in\{0,1\}$ is easy to reach a contradiction. Then t=1+3kt=-1+3k substituting again
$t^{2}-t+1$ $=$ $9k^{2}-9k+3,$
$t^{3}-6t^{2}+3t+1$ $=$ $27k^{3}-81k^{2}+54k-9.$
Therefore, $v_{3}(t^{3}-6t^{2}+3t+1)=2$ and $v_{3}(t^{2}-t+1)=1$. So (22) becomes
@@
If $E(\mathbb{Q})[3]=\mathcal{C}_{3}$, the hypothesis implies that $E_{\mathrm{tors}}(K)=E_{\mathrm{tors}}(\mathbb{Q})\times\mathcal{C}_{3}$, because of Theorem 1. Therefore we can assume that $E(\mathbb{Q})[3]$ is trivial.
@@
Now we go by contradiction. Using Theorem 1, we see that the possible cases that can arise are:
@@
  • $E_{\mathrm{tors}}(\mathbb{Q})=\mathcal{C}_{4}$ and $E_{\mathrm{tors}}(K)=\mathcal{C}_{2}\times\mathcal{C}_{12}$,
@@
Case I: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$ and $\mathcal{C}_{4}\leq E_{\mathrm{tors}}(K)$. As in subsection 3 we can do a change of variables and we get a Weierstrass equation:
@@
Now Lemma 1 tell us that $K=\mathbb{Q}(\sqrt{A\pm 2s})$ with $s\in\mathbb{Q}$ such that $s^{2}=B$, so $s\in\mathbb{Z}$. As $3|d$, $A\pm 2s$ must be a multiple of $3$. Then, from the equation relating the discriminants, we obtain
@@
Case II: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}$ and $\mathcal{C}_{2}\times\mathcal{C}_{2}\leq E_{\mathrm{tors}}(K)$. We can compute, as in Section 2, the $2$-torsion points with the $2$-division polynomial $\psi_{2}=2y+a_{1}x+a_{3}$; substituting
@@
we obtain the expression (see, again [14, Ch. III])
@@
$\Delta=4^{2}(\alpha-\beta)^{2}(\alpha-c)^{2}(\beta-c)^{2}\;\Longrightarrow\;0=2v_{3}(\alpha-\beta)+2v_{3}(\alpha-c)+2v_{3}(\beta-c).$
@@
Case III: $E_{\mathrm{tors}}(\mathbb{Q})[2]=\mathcal{C}_{2}\times\mathcal{C}_{2}$ and $\mathcal{C}_{4}\times\mathcal{C}_{2}\leq E_{\mathrm{tors}}(K)$. Again, as in Section 3, we can do a change of variables and we get a Weierstrass equation:
@@
In this situation, like in Section 3, we can assume that there is a point $Q\in E(K)[4]$ such that $2Q=(0,0)$. Applying Lemma 2 as in Subsection 3, we can assume that the polynomial $x^{3}+Ax^{2}+Bx$ factors as $x(x-\alpha)(x-\beta)$, where $\alpha,\beta\in\mathbb{Q}$ satisfy that $\alpha$ and $\beta$ are squares in $K^{*}$, but they are not both squares in $\mathbb{Q}^{*}$. Thus either $\alpha=a^{2}$ and $\beta=db^{2}$, or $\alpha=da^{2}$ and $\beta=b^{2}$, or else $\alpha=da^{2}$ and $\beta=db^{2}$ for some rational integers $a$ and $b$. Replacing $B=\alpha\beta$ and $A=-\alpha-\beta$ in the equation for $\Delta$ above, we obtain in all cases that there is an integer $V\in\mathbb{Z}$ such that
@@ 5. The primes $\ell=5$ and $\ell=7$
Throught this section we can already assume that the curve $E$ does not have complex multiplication, as the CM case has already been solved by E. González–Jiménez [2]. Our goal here is to prove Theorem 3, as stated in the introduction.
Theorem 3. Let $E/\mathbb{Q}$ be an elliptic curve, $K$ a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$. If $p=5,7$ ramifies in $K$, then $p|N_{E}$.
@@
Next, we will see the form of $I(\Lambda/\ell)$ according to the type of reduction of $E$. The following propositions, which are proved in [13, Section 1.11, Section 1.12], exactly determine the image of $I_{\ell}$ under $\overline{\rho}_{E,\ell}$:
Proposition 9.
Let $E/\mathbb{Q}$ be an elliptic curve with good reduction of height 1 or multiplicative reduction in a prime $\ell$. Then one and only one of the following possibilities holds:
@@
    The wild inertia group $I^{\mathrm{wild}}_{\ell}$ does not act trivially on $E[\ell]$. Then the image of $I^{\mathrm{wild}}_{\ell}$ under $\overline{\rho}_{E,\ell}$ is a cyclic group of order $\ell$, and in a suitable basis, it can be represented as
@@
Proposition 10.
Let $p,\ell$ be two different primes. Let $E/\mathbb{Q}$ be an elliptic curve with multiplicative reduction in a prime $p$. Then the image of $I_{\ell}$ is trivial or a cyclic group of order $\ell$.
We will now analyze each of the possible types of reduction of $E$ at $\ell$, proving that, for each case, the existence of a point $P\in E(K)[\ell]\setminus E(\mathbb{Q})$ yields a contradiction. Taking Remark 1 into account, this reasoning will prove Theorem 3.
5.1. Ordinary good reduction or multiplicative reduction, with nontrivial wild inertia action
We will assume in this case that $E$ has good reduction at $\ell$ and $I^{\mathrm{wild}}_{\ell}$ does not act trivially. Note that $H_{3}$ does not leave any element other than $(0,0)^{t}$ invariant; in particular, there cannot be any rational point of $\ell$-torsion.
Furthermore, no subgroup of index $2$ in $H_{3}$ leaves any element other than $(0,0)^{t}$ invariant. Indeed, since the cardinality of $H_{3}$ is $\ell(\ell-1)$ and $\ell>2$, any subgroup $H$ of $H_{3}$ of index $2$ must contain an element of order $\ell$. Such elements are precisely the non-trivial elements of $H_{2}$. Therefore, $H$ contains $H_{2}$. On the other hand, we have
@@ 5.1. Ordinary good reduction or multiplicative reduction, with nontrivial wild inertia action
if and only if $by=0$, so the only nontrivial fixed points by the entire subgroup are of the form $(x,0)^{t}$, where $x\neq 0$. However, any subgroup of index $2$ must contain at least one matrix of the form
@@
(otherwise the subgroup would be contained in $H_{2}$, and $H_{2}$ is already too small to have index two in $H_{3}$, provided $\ell>3$). Now, this matrix does not fix the point $(x,0)^{t}$. This shows that for $\ell=5,7$, it is not possible to add an $\ell$-torsion point over a quadratic extension that ramifies at $\ell$.
5.2. Ordinary good reduction or multiplicative reduction, with trivial wild inertia action
@@
Case 5.2.I: $\ell=5$. In this case, [17] shows that $\overline{\rho}_{E,5}(G_{\mathbb{Q}})$ is conjugate in $\mathop{\mathrm{GL}}\nolimits_{2}(\mathbb{F}_{5})$ to a group from a list of 15 possible groups (see [17, Theorem 1.4]). Furthermore, Sutherland studies these groups and determines in each case the index of the largest subgroup that fixes a nonzero vector in $\mathbb{F}_{5}^{2}$; this quantity coincides with the degree of the minimal extension $K/\mathbb{Q}$ such that $E$ has a rational point of $5$-torsion.
In [15, Table 3, p. 64], we can find the list of these 15 groups along with their indices. There are only four of them where this degree is $2$, specifically those labeled as 5Cs.1.3, 5Cs.4.1, 5B.1.4, and 5B.4.1. Let us examine each of these cases:
  • 5Cs.1.3. It is a cyclic subgroup of order $4$, generated by the matrix
    $\begin{pmatrix}3&0\\ 0&4\end{pmatrix}$.
    Since we are assuming that $E$ has good ordinary reduction at $5$, $\overline{\rho}_{E,\ell}(G_{\mathbb{Q}})$ must contain a subgroup conjugate to $H_{1}$, which is also cyclic of order $4$. However, the subgroup 5Cs.1.3 is not conjugate to $H_{1}$, because it does not fix any nonzero element of $\mathbb{F}_{5}^{2}$. The conclusion is that this image cannot occur if $E$ has good ordinary reduction at $5$ (it also cannot have supersingular reduction, by a similar reasoning).
@@
    The subgroup generated by the matrix
    $\begin{pmatrix}1&0\\ 0&2\end{pmatrix}$
    is conjugate to the subgroup $H_{1}$, and it is the only cyclic subgroup of order $4$ in 5Cs.4.1 that fixes a nonzero element of $\mathbb{F}_{5}^{2}$. Therefore,
@@
    Moreover, $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$ is a subgroup of $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/\mathbb{Q})$ of order $4$. Since $L\neq K$, it must be a subgroup of $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/\mathbb{Q})$ different from $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)$, and thus it cannot fix a nonzero element of $\mathbb{F}_{5}^{2}$.
@@
    Again, there is a subgroup which is conjugate to $H_{1}$, the subgroup
    $H=\left\{\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$.
@@
    In this case there are more cyclic subgroups of order $4$ that have a fixed point; specifically:
    $\left\{\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\begin{pmatrix}1&3\\ 0&4\end{pmatrix},\begin{pmatrix}1&2\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\},\quad\left\{\begin{pmatrix}1&1\\ 0&3\end{pmatrix},\begin{pmatrix}1&4\\ 0&4\end{pmatrix},\begin{pmatrix}1&3\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&3\end{pmatrix}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\},\quad\left\{\begin{pmatrix}1&3\\ 0&3\end{pmatrix},\begin{pmatrix}1&2\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$.
    In fact, all of these groups are conjugate to $H$ inside the group $5B.4.1$:
    $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\begin{pmatrix}1&1\\ 0&1\end{pmatrix}^{-1}=\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\quad\begin{pmatrix}1&1\\ 0&2\end{pmatrix}\begin{pmatrix}1&0\\ 0&3\end{pmatrix}\begin{pmatrix}1&1\\ 0&2\end{pmatrix}^{-1}=\begin{pmatrix}1&1\\ 0&3\end{pmatrix}$,
    $\begin{pmatrix}1&2\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\begin{pmatrix}1&2\\ 0&1\end{pmatrix}^{-1}=\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\quad\begin{pmatrix}1&3\\ 0&2\end{pmatrix}\begin{pmatrix}1&0\\ 0&3\end{pmatrix}\begin{pmatrix}1&3\\ 0&2\end{pmatrix}^{-1}=\begin{pmatrix}1&3\\ 0&3\end{pmatrix}$.
@@
    By choosing a basis, we can assume that $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)$ is one of them. If there exists a point of $5$-torsion over $K$, the only possibility is that $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$, which is a group of order $20$, is the union of all elements whose upper left entry is $1$. This can be seen by looking at the list of elements of the group 5B.4.1; the only 20 elements that fix the same element of $\mathbb{F}_{5}^{2}$ are these.
@@
    But in that case, $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/L)\subset\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[5])/K)$, and therefore $K\subset L$. Since we have seen that $K$ and $L$ are linearly disjoint over $\mathbb{Q}$, this is a contradiction.